OpenStudy (anonymous):

at the given point fine the slope of the curve the line that is tangent to the curve or the line that is normal to the curve. (2x^2)y - pi cos y =3pi, tangent at (1,pi)

OpenStudy (anonymous):

?

sam (.sam.):

"fine" or "find"? lol

OpenStudy (anonymous):

maybe at the given point find the slope of the line that is tangent to the curve and the line that is normal to the curve

OpenStudy (anonymous):

curve being \[2x^2y-\pi\cos(y)=3\pi\]?

OpenStudy (hihi67):

@satellite73 (a^2)^2/3b / 4a^5/(b^3)^2=

OpenStudy (anonymous):

something wrong with this problem, since if you replace x by 1 i don't think you get \[\pi\]

OpenStudy (anonymous):

Yea i meant find, my bad

OpenStudy (anonymous):

2x2y−πcos(y)=3π This is the equation and there asking me to find the tangent line at (1, pi)